# LawChain: Modeling Legal Reasoning Chains for Chinese Tort Case Analysis

### v1

@@ Abstract
Legal reasoning is a fundamental component of legal analysis and decision-making. Existing computational approaches to legal reasoning predominantly rely on generic reasoning frameworks such as syllogism and IRAC, which do not comprehensively examine the nuanced processes that underpin legal reasoning. Moreover, current research has largely focused on criminal cases, with insufficient modeling for civil cases. In this work, we present a novel framework for explicitly modeling legal reasoning in the analysis of Chinese tort-related civil cases. We first operationalize the legal reasoning processes used in tort analysis into the LawChain framework. LawChain is a three-module reasoning framework, with each module consisting of multiple finer-grained sub-steps. Informed by the LawChain framework, we introduce the task of tort legal reasoning and construct an evaluation benchmark, LawChaineval, to systematically assess the critical steps within analytical reasoning chains for tort analysis. Leveraging this benchmark, we evaluate state-of-the-art large language models for their legal reasoning ability in civil tort contexts. Our results indicate that current models still fall short in accurately handling crucial elements of tort legal reasoning. Furthermore, we introduce several baseline approaches that explicitly incorporate LawChain-style reasoning through prompting or post-training. We conduct further experiments on additional legal analysis tasks, such as Legal Named-Entity Recognition and Criminal Damages Calculation, to verify the generalizability of these baselines. The proposed baseline approaches achieve significant improvements in tort-related legal reasoning and generalize well to related legal analysis tasks, thus demonstrating the value of explicitly modeling legal reasoning chains to enhance the reasoning capabilities of language models.
@@ Introduction
The rapid development of large language models (LLMs) has led to widespread adoption of these tools across diverse domains. Recent models, such as Gemini-2.5-Pro (Comanici et al. 2025) and Qwen-3 (Yang et al. 2025), now position reasoning as a default and essential component of their architecture. While notable progress has been made in advancing the reasoning abilities of LLMs in areas such as mathematics and coding (Wei et al. 2022; Huang et al. 2023; El-Kishky et al. 2025), LLMs’ reasoning capabilities in the legal domain, where nuanced outputs are dependent on structured and domain-specific legal reasoning, remain underexplored.
Legal reasoning is a foundational element of legal analysis and decision-making. Legal scholars argue that there is, descriptively, something unique that can be characterized as “thinking like a lawyer” (Samuelson 1997; Schauer 2009). This encompasses distinctive forms of reasoning, such as the systematic application of codified legal rules, the decomposition of legal questions into granular elements for structured judgment, and ensuring that similar cases are treated uniformly. Better modeling of legal reasoning processes not only improves the accuracy and efficiency of legal analysis, but also enhances the interpretability of results by providing explicit reasoning chains.
Figure 1: Illustration of the case information (English translation) in a civil tort case and the three-module LawChain reasoning framework.
However, existing work on computational legal reasoning has primarily focused on generic legal reasoning frameworks, such as syllogism (Deng et al. 2023; Jiang and Yang 2023) and IRAC (Yu, Quartey, and Schilder 2023; Kuppa, Rasumov-Rahe, and Voses 2023; Servantez et al. 2024), often simulating reasoning through simple prompting strategies. In addition, research has concentrated predominantly on criminal law, with insufficient attention paid to civil cases, especially tort cases, which are highly relevant to everyday life. Tort law plays a fundamental role in modern legal systems, in both protecting individual rights and compensating plaintiffs for harms suffered.
In this work, we explicitly model legal reasoning chains for tort-related civil cases within the context of Chinese civil law. Drawing on established legal theory (Smith 1911; Moore 1999), we operationalize the analytical process of tort adjudication into the LawChain reasoning framework - a multi-module reasoning chain in which each module comprises a set of key steps or elements essential to tort legal reasoning (see Figure 1 for an illustration).
@@
Leveraging our LawChain reasoning framework, we further implement reasoning-enhanced baselines using approaches such as legal prompting, supervised fine-tuning (SFT) and direct preference optimization (DPO). Results demonstrate that integrating explicit legal reasoning chains through these techniques consistently improves model performance on identifying key elements for tort case analysis, underscoring the value of explicit legal reasoning modeling in enhancing LLMs’ legal reasoning abilities.
Finally, we evaluate our fine-tuned baselines on related legal AI tasks, including Legal Named-Entity Recognition and Criminal Damages Calculation (Fei et al. 2024). Our reasoning-enhanced models achieve comparable or superior performance to zero-shot counterparts on these tasks, demonstrating the generalizability of reasoning-enhanced approaches to other legal AI tasks.
@@
  • Building on this framework, we introduce the novel task of tort legal reasoning and construct a dedicated benchmark to evaluate models’ performance on this task. To the best of our knowledge, the constructed benchmark is the first evaluation dataset that focuses on the assessment of legal reasoning abilities in the context of Chinese civil tort law. In addition, we curate a reasoning-enhanced training dataset to enable effective learning of legal reasoning patterns.
  • Our empirical results demonstrate that current LLMs exhibit notable limitations in tort-related legal reasoning, particularly in labor dispute cases. Furthermore, we find that generic syllogistic reasoning frameworks do not yield meaningful improvements in this domain, underscoring the need to move beyond generic approaches and adopt domain-specific reasoning paradigms tailored to the complexity of civil tort law. We further show that incorporating LawChain-style reasoning through prompting, SFT and DPO leads to substantial performance gains on the tort reasoning task. These improvements also generalize effectively to other related tasks, underscoring the value of explicitly modeling structured legal reasoning in enhancing LLM performance on legal analysis tasks.
@@
Legal reasoning refers to the analytical process by which legal practitioners integrate legal norms with case facts to reach a legal conclusion (Samuelson 1997; Schauer 2009). Legal reasoning holds a central role in judicial practice, serving as the foundation for legal analysis and as a mechanism to ensure the consistency of judicial decisions. Distinct from general forms of logical reasoning, legal reasoning is shaped by the constraints of legal systems, precedential rules and various institutional factors (Scharffs 2004). The diversity and complexity of real-world cases also make the practical application of legal reasoning highly challenging.
Tort liability is a form of civil liability that aims to remedy infringed civil rights and safeguard the legitimate interests of civil subjects (Smith 1911). In judicial practice, tort disputes are of particular significance, as they encompass a broad spectrum of harms, ranging from personal injury and property damage to emotional distress and online harassment, permeating many aspects of everyday life (Latin 1985; Wagner 2019). In comparison to other areas of law, tort liability cases exhibit distinctive patterns in their reasoning pathways. A foundational step in the analysis of tort disputes is the determination of liability, which involves a systematic examination of the constituent elements: conduct, harm, causation and fault (Smith 1911; Catala and Weir 1965; Moore 1999). Once liability is established, legal practitioners must further consider the nature of the tortious act and the severity of the harm to determine the appropriate forms of liability, with monetary compensation being the most prevalent remedy (Van Wijck and Winters 2001). In certain cases, additional complexities such as compensation for emotional distress and the application of rules for loss offset and comparative fault may arise (Pearson 1979; Kontorovich 2001), demanding commensurately complex legal reasoning to successfully adjudicate.
@@
Computational modeling of legal reasoning aims to simulate the multi-step inferential processes underpinning legal decision-making (Deng et al. 2023; Fernandes et al. 2025). In recent years, structured prompting techniques, particularly those that extend chain-of-thought (CoT) prompting (Wei et al. 2022) to legal tasks (Yu, Quartey, and Schilder 2022; Jiang and Yang 2023), have garnered increasing prominence. A significant line of work leverages syllogism-based prompting (Jiang and Yang 2023), which encourages models to organize their reasoning in a triadic structure: major premise (legal norm), minor premise (case facts) and conclusion (judgment). Other studies have incorporated schema-based reasoning frameworks such as IRAC (Yu, Quartey, and Schilder 2023; Kuppa, Rasumov-Rahe, and Voses 2023), encouraging generated outputs to resemble authentic legal argumentation. Another logic-based approach (Servantez et al. 2024) focuses on atomistic parsing of legal rules, decomposing complex provisions into atomic Boolean conditions, allowing for fine-grained alignment with statutory interpretation practices.
In addition to prompting methods, supervised fine-tuning has been used to integrate legal reasoning patterns into models. For instance, syllogism-based fine-tuning (Deng et al. 2023) encodes syllogistic legal deduction into models, reinforcing deductive reasoning during inference. Similarly, schema-based fine-tuning (Fernandes et al. 2025) trains models to transform case analyses into structured IRAC representations, effectively internalizing schema-based scaffolds within models’ reasoning process.
@@
LawChain: Benchmarking Legal Reasoning
To model legal reasoning in tort law, we first conceptualize tort-related decision-making as a structured, multi-step reasoning process, which we further operationalize as a legal reasoning chain, referred to here as LawChain. Stemming from this conceptual base, we propose a novel task of tort legal reasoning and construct an evaluation dataset (LawChaineval). We further construct a reasoning-enhanced training dataset (LawChaintrain) that incorporates explicit LawChain-style reasoning tracing, which can be used to fine-tune models for improved legal reasoning.
@@
Tort law encompasses a broad spectrum of legally protected rights and interests, and thus is relevant to diverse aspects of everyday life. Tort cases exhibit distinctive reasoning structures, particularly centered around the establishment and apportionment of liability. The analytical process typically begins with determining whether liability exists, and proceeds to assess the severity and consequences of the tortious act in order to determine the extent of legal responsibility and appropriate forms of compensation.
Drawing on established judicial reasoning practices, we operationalize tort analysis into a structured, three-module reasoning chain. Each module is further decomposed into a list of specific legal sub-tasks, operationalizing the abstract process of legal reasoning into concrete, verifiable components suitable for computational modeling and evaluation.
The LawChain reasoning framework consists of three modules: (1) a legal element identification module, which identifies relevant foundational information necessary for legal analysis, (2) a liability analysis module, the core module, which centers on determining and apportioning liability, and (3) a judgment summarization module, which synthesizes the outcomes of the preceding modules into a final legal decision. An illustration of the LawChain reasoning framework is presented in Figure 1.
@@
Liability analysis module. As the central component of the LawChain, this module determines whether the defendant(s) should bear legal liability, and if so, how such liability should be apportioned. The reasoning process is divided into two sub-tasks for each defendant:
  • Liability determination. This sub-task assesses whether the defendant has committed a legally recognized tort and thus should be held civilly liable. Following legal doctrine, we structure this analysis around four foundational elements of tort liability: conduct, harm, causation and fault (Smith 1911). The presence and interaction of these elements are evaluated to determine both the existence and extent (i.e., proportional liability) of a defendant’s responsibility.
  • Liability apportionment. Once liability is established, this sub-task determines the specific forms of legal responsibility to be imposed, in accordance with civil law. Common forms include monetary compensation, the restitution of damages or cessation of infringement (Wan et al. 2022). Given that monetary compensation is the predominant remedy in tort law (Van Wijck and Winters 2001), we explicitly add a step of damages calculation where the total loss suffered by the victim is estimated. The total of estimated damages is multiplied by the proportion of liability attributed to each defendant to obtain a precise amount of compensation. Additional remedies, such as restitution or cessation of infringement, are also considered where legally applicable.
@@
Facilitated by the modular and stepwise structure formalized in the LawChain reasoning framework, we introduce a novel task of tort legal reasoning (TLR) to systematically evaluate LLMs’ capacity to perform structured legal reasoning in the context of tort law. We construct an evaluation benchmark, LawChaineval, to assess models’ ability to identify and reason through the critical components involved in tort case analysis. In addition, we propose an LLM-assisted scoring framework to quantify model performance at each reasoning step, enabling interpretable and modular evaluation.
@@
Tort legal reasoning (TLR) requires a model to correctly identify or predict a structured sequence of key legal elements that together form a coherent reasoning chain for tort adjudication. The LawChain framework provides a principled decomposition of the reasoning process into discrete modules, each associated with legally meaningful sub-tasks. We construct the TLR task by closely aligning the task with this formalization, allowing models to be evaluated across the entire reasoning trajectory, from case setup to judgment formulation.
@@
We construct the evaluation dataset LawChaineval based on real-world judicial documents. Authentic judicial decisions offer a realistic, diverse source of tort-related legal scenarios. In particular, cases that proceed to litigation typically involve greater factual and legal complexity, as they often center on issues of conceptual ambiguity or interpretive dispute (Schauer 2009). These characteristics make litigated cases especially well-suited for modeling and evaluating tort legal reasoning.
We start by sampling 1,000 real-world tort case judgments from China Judgments Online (2013). To ensure that the evaluation dataset centers on substantive legal reasoning, we exclude cases involving procedural matters, such as voluntary withdrawals, jurisdictional transfers or dismissals on procedural grounds during the sampling process. The resulting dataset comprehensively covers major types of tort disputes, encompassing 45 distinct tort subcategories.
We design an LLM-assisted extraction pipeline to systematically retrieve key legal information from court decision texts. This pipeline extracts essential elements required for tort reasoning, including: plaintiff(s), defendant(s), type of dispute, applicable legal statutes, claims of the plaintiff(s), case facts, total damages, liability determinations and judgment outcomes. For elements commonly expressed in structured formats in court decisions, such as parties, claims, citations of law and judgments, we employ rule-based extraction techniques, including keyword matching and regular expression-based heuristics. For elements requiring holistic understanding and document-wise summarization, such as fact descriptions, liability reasoning and damages calculation, we utilize DeepSeek-V3 (Liu et al. 2024) to synthesize relevant information from the court decision texts.
To ensure the quality of data present in LawChaineval, we conduct a rigorous manual validation process on the automatically extracted case information. All extracted data are reviewed and refined by a trained legal annotator, under the supervision of a senior researcher with extensive experience in legal AI and annotation workflows. The extraction and manual validation processes ensure a high-quality benchmark for evaluating tort-related legal reasoning.
@@
To systematically evaluate model outputs on the tort legal reasoning task, we adopt an “LLM-as-a-Judge” evaluation paradigm, widely used in recent literature due to its demonstrated effectiveness and scalability for automated evaluation (Zheng et al. 2023; Gu et al. 2024). We design a structured and automated scoring process using GPT-4o (Hurst et al. 2024) as the evaluation model. Importantly, instead of relying on subjective, holistic assessments, we decompose the evaluation into a series of discrete, well-defined sub-dimensions and prompt GPT-4o to make rule-based, factual judgments for each sub-dimension. This decomposition mitigates the risk of evaluation bias and enhances reliability of LLM-based evaluation, aligning with existing practices in recent studies (Liu et al. 2023; Wang et al. 2024; Gu et al. 2024).
The evaluation covers seven distinct scoring dimensions, corresponding directly to the core reasoning elements examined in the LawChaineval benchmark. For each test instance, GPT-4o is instructed to evaluate the model’s response against gold-standard answers, i.e., reference case elements provided in the LawChaineval dataset. The scoring schema for each reasoning element is defined as follows:
@@
    Liability apportionment: Measures the model’s ability to identify the liable parties and to correctly determine the proportion of liability, forms of civil responsibility (e.g., compensation, restitution) and the associated factual and legal justifications.
@@
To assess the robustness and reliability of the LLM-based evaluation, we conduct a quality validation procedure where we randomly sample 300 instances from GPT-4o’s ratings and ask two independent annotators to review the ratings across seven scoring dimensions. For each of GPT-4o’s ratings, annotators independently assign a binary judgment: a score of 1 if they consider GPT-4o’s evaluation correct, and 0 otherwise. In cases where the two annotators disagree, a senior researcher with expertise in legal AI practices is consulted, and the annotators engage in discussion to reach a consensus. The final agreed-upon labels (0 or 1) are used as the reference for calculating the accuracy of GPT-4o’s ratings. All initial disagreements are also recorded and incorporated into the calculation of the inter-annotator agreement score to ensure a fair assessment. Based on this validation criterion, GPT-4o achieves an overall rating accuracy of 0.949 on the sampled set. Inter-annotator agreement, measured using Cohen’s Kappa (Cohen 1960), yields a score of 0.619, indicating relatively strong consistency between annotators [^1] According to Landis and Koch (1977), a Cohen’s Kappa score between 0.61 and 0.80 signifies “substantial agreement”..
Constructing LawChain-Style Training Data
@@
To prepare training data enriched with legal reasoning information, we sample approximately 10,600 court decision documents for tort cases. For each case, we first extract key legal elements following the same information extraction practice described in LawChaineval construction. We then leverage DeepSeek-V3 (Liu et al. 2024) to assist in the automatic construction of legal reasoning chains, building upon the extracted legal elements. For each step in the reasoning process, we determine the specific types of information required to infer towards the current reasoning objective and construct corresponding reasoning formulas. For example, the identification of litigation parties (i.e., plaintiff(s) and defendant(s)) requires consideration of both the case facts and the claims of the plaintiff(s), while the calculation of compensation details depends on the previously determined liability ratio and the victim’s actual losses. An illustrative example is presented in Figure 2. Each constructed reasoning chain is enclosed in a <<think>> and <</think>> tag pair and prefixed to the final answer consisting of the extracted elements, resulting in the LawChaintrain training data.
@@
We conduct human validation to assess the quality of the reasoning chains automatically generated by DeepSeek-V3. Two annotators independently evaluate 300 samples of LLM-generated reasoning chains based on two assessment criteria: consistency and coherence. Consistency measures the alignment of the generated reasoning text with the original legal elements from which it is constructed, while coherence evaluates whether the output forms a logically sound and well-structured reasoning narrative. The training data constructed by DeepSeek-V3 obtains a consistency rate of 0.973 and a coherence rate of 0.993. The inter-annotator agreement between the two annotators, measured as the Cohen’s Kappa score (Cohen 1960), is 0.866, indicating a strong agreement between annotators.
@@
Table 1: Evaluation results on the LawChaineval benchmark. Entries labeled with the model name alone (e.g., GPT-4o) represent the zero-shot inference setting. The variant “w/ PromptLC” denotes models prompted using LawChain-style legal reasoning prompts. Variants “w/ SFTSyll” and “w/ SFTLC” denote models fine-tuned on syllogism-style and LawChain-style SFT data, respectively. The variant “w/ DPOLC” refers to models optimized with LawChain-style DPO data.

### v2

# LexChain: Modeling Legal Reasoning Chains for Chinese Tort Case Analysis

@@ Abstract
Legal reasoning is a fundamental component of legal analysis and decision-making. Existing computational approaches to legal reasoning predominantly rely on generic reasoning frameworks such as syllogism, which do not comprehensively examine the nuanced process of legal reasoning. Moreover, current research has largely focused on criminal cases, with insufficient modeling for civil cases. In this work, we present a novel framework to explicitly model legal reasoning in the analysis of Chinese tort-related civil cases. We first operationalize the legal reasoning process in tort analysis into the three-module LexChain framework, with each module consisting of multiple finer-grained sub-steps. Informed by the LexChain framework, we introduce the task of tort legal reasoning and construct an evaluation benchmark to systematically assess the critical steps within analytical reasoning chains for tort analysis. Leveraging this benchmark, we evaluate existing large language models for their legal reasoning ability in civil tort contexts. Our results indicate that current models still fall short in accurately handling crucial elements of tort legal reasoning. Furthermore, we introduce several baseline approaches that explicitly incorporate LexChain-style reasoning through prompting or post-training. The proposed baselines achieve significant improvements in tort-related legal reasoning and generalize well to related legal analysis tasks, demonstrating the value of explicitly modeling legal reasoning chains to enhance the reasoning capabilities of language models.
Data and codehttps://github.com/thunlp/LexChain
@@ Introduction
Figure 1: Illustration of the case information (English translation) in a Chinese tort case and the three-module LexChain reasoning framework.
The rapid development of large language models (LLMs) has led to widespread adoption of these tools across diverse domains. Recent models, such as Gemini-2.5-Pro (comanici2025gemini) and Qwen-3 (yang2025qwen_3), now position reasoning as a default and essential component of their architecture. While notable progress has been made in advancing the reasoning abilities of LLMs in areas such as mathematics and coding (wei2022chain; huang2023codecot; el2025competitive), LLMs’ reasoning capabilities in the legal domain, where nuanced outputs are dependent on structured and domain-specific legal reasoning, remain underexplored.
Legal reasoning is a foundational element of legal analysis and decision-making. Legal scholars argue that there is, descriptively, something unique that can be characterized as “thinking like a lawyer” (samuelson1997introducing; schauer2009thinking). This encompasses distinctive forms of reasoning, such as the systematic application of codified legal rules, the decomposition of legal questions into granular elements for structured judgment, and ensuring that similar cases are treated uniformly. Better modeling of legal reasoning not only improves the accuracy and efficiency of legal analysis, but also enhances the interpretability of results by providing explicit reasoning chains.
However, existing work on computational legal reasoning has primarily focused on generic legal reasoning frameworks, such as syllogism (deng2023syllogistic; jiang2023legal) and IRAC (yu2023exploring; kuppa2023chain; servantez2024chain), often simulating reasoning through simple prompting strategies. In addition, research has concentrated predominantly on criminal law, with insufficient attention paid to civil cases, especially tort cases, which are highly relevant to everyday life. Tort law plays a fundamental role in modern legal systems, in both protecting individual rights and compensating plaintiffs for harms suffered.
In this work, we explicitly model legal reasoning chains for tort-related civil cases within the context of Chinese civil law. Drawing on established legal theory (smith1911legal; moore1999causation), we operationalize the analytical process of tort adjudication into the LexChain reasoning framework—a multi-module reasoning chain in which each module comprises a set of key steps or elements essential to tort legal reasoning (see Figure 1 for an illustration).
@@
Leveraging our LexChain reasoning framework, we further implement reasoning-enhanced baselines using approaches such as legal prompting, supervised fine-tuning (SFT) and direct preference optimization (DPO). Results demonstrate that integrating explicit legal reasoning chains through these techniques consistently improves model performance on identifying key elements for tort case analysis, underscoring the value of explicit legal reasoning modeling in enhancing LLMs’ legal reasoning abilities.
Finally, we evaluate our fine-tuned baselines on related legal AI tasks, including Legal Named-Entity Recognition and Criminal Damages Calculation (fei2024lawbench). Our reasoning-enhanced models achieve comparable or superior performance to zero-shot counterparts on these tasks, demonstrating the generalizability of reasoning-enhanced approaches to other legal AI tasks.
@@
  • We introduce the novel task of tort legal reasoning and construct a dedicated benchmark to evaluate models’ performance on this task. To the best of our knowledge, the constructed benchmark is the first evaluation dataset that focuses on the assessment of legal reasoning abilities in the context of Chinese civil tort law. In addition, we curate a reasoning-enhanced training dataset to enable effective learning of legal reasoning patterns.
  • Our empirical results demonstrate that current LLMs exhibit notable limitations in tort-related legal reasoning, particularly in labor dispute cases. Furthermore, we find that generic syllogistic reasoning frameworks do not yield meaningful improvements in this domain, underscoring the need to move beyond generic approaches and adopt domain-specific reasoning paradigms tailored to the complexity of civil tort law. We further show that incorporating LexChain-style reasoning through prompting, SFT and DPO leads to substantial performance gains on the tort reasoning task. These improvements also generalize effectively to other related tasks, underscoring the value of explicitly modeling structured legal reasoning in enhancing LLM performance on legal analysis tasks.
@@
Legal reasoning refers to the analytical process by which legal practitioners integrate legal norms with case facts to reach a legal conclusion (samuelson1997introducing; schauer2009thinking). Legal reasoning holds a central role in judicial practice, serving as the foundation for legal analysis and as a mechanism to ensure the consistency of judicial decisions. Distinct from general forms of logical reasoning, legal reasoning is shaped by the constraints of legal systems, precedential rules and various institutional factors (scharffs2004character). The diversity and complexity of real-world cases also make the practical application of legal reasoning highly challenging.
Tort liability is a form of civil liability that aims to remedy infringed civil rights and safeguard the legitimate interests of civil subjects (smith1911legal). In judicial practice, tort disputes are of particular significance, as they encompass a broad spectrum of harms, such as personal injury, property damage and online harassment, permeating many aspects of everyday life (latin1985problem; wagner2006comparative). In comparison to other areas of law, tort liability cases exhibit distinctive patterns in their reasoning pathways. A foundational step in tort analysis is the determination of liability, which involves a systematic examination of the constituent elements: conduct, harm, causation and fault (smith1911legal; catala1965delict; moore1999causation). Once liability is established, legal practitioners must further consider the nature of the tortious act and the severity of the harm to determine the appropriate forms of liability, with monetary compensation being the most prevalent remedy (van2001principle). In certain cases, additional complexities such as compensation for emotional distress and the application of rules for loss offset may arise (pearson1979apportionment; kontorovich2001mitigation), demanding complex legal reasoning to successfully adjudicate.
@@
Computational modeling of legal reasoning aims to simulate the multi-step inferential processes underpinning legal decision-making (deng2023syllogistic; fernandes2025llama). In recent years, structured prompting techniques (yu2022legal; jiang2023legal) have garnered increasing prominence. A significant line of work leverages syllogism-based prompting (jiang2023legal), which encourages models to organize their reasoning in a triadic structure: major premise (legal norm), minor premise (case facts) and conclusion (judgment). Other studies have incorporated schema-based reasoning frameworks such as IRAC (yu2023exploring; kuppa2023chain), encouraging generated outputs to resemble authentic legal argumentation. Another logic-based approach (servantez2024chain) focuses on atomistic parsing of legal rules, decomposing complex provisions into atomic Boolean conditions, allowing for fine-grained alignment with statutory interpretation practices.
In addition to prompting methods, supervised fine-tuning has been used to integrate legal reasoning patterns into models. For instance, syllogism-based fine-tuning (deng2023syllogistic) encodes syllogistic legal deduction into models, reinforcing deductive reasoning during inference. Similarly, schema-based fine-tuning (fernandes2025llama) trains models to transform case analyses into structured IRAC representations, effectively internalizing schema-based scaffolds within models’ reasoning process.
@@
LexChain: Benchmarking Legal Reasoning
To model legal reasoning in tort law, we first conceptualize tort-related decision-making as a structured, multi-step reasoning process, which we further operationalize as a legal reasoning chain, referred to here as LexChain. Stemming from this conceptual base, we propose a novel task of tort legal reasoning and construct an evaluation dataset (LexChaineval). We further construct a reasoning-enhanced training dataset (LexChaintrain) that incorporates explicit LexChain-style reasoning tracing, which can be used to fine-tune models for improved legal reasoning.
@@
Tort analysis exhibits distinctive reasoning structures, particularly centered around the establishment and apportionment of liability. The analytical process typically begins with determining whether liability exists, and proceeds to assess the severity and consequences of the tortious act in order to determine the extent of legal responsibility and appropriate forms of compensation.
Drawing on established legal frameworks (smith1911legal; catala1965delict; van2001principle) and consultations with practicing judges and lawyers, we formalize tort analysis into a structured, three-module reasoning chain. Each module is further decomposed into a list of specific legal sub-tasks, operationalizing the abstract process of legal reasoning into concrete, verifiable components suitable for computational modeling and evaluation.
The LexChain reasoning framework consists of three modules: (1) a legal element identification module, which identifies relevant foundational information necessary for legal analysis, (2) a liability analysis module, the core module, which centers on determining and apportioning liability, and (3) a judgment summarization module, which synthesizes the outcomes of the preceding modules into a final legal decision. An illustration of the LexChain reasoning framework is presented in Figure 1.
@@
Liability analysis module. As the central component of the LexChain, this module determines whether the defendant(s) should bear legal liability, and if so, how such liability should be apportioned. The reasoning process is divided into two sub-tasks for each defendant:
  • Liability determination. This sub-task assesses whether the defendant has committed a legally recognized tort and thus should be held civilly liable. Following legal doctrine, we structure this analysis around four foundational elements of tort liability: conduct, harm, causation and fault (smith1911legal). The presence and interaction of these elements are evaluated to determine both the existence and extent (i.e., proportional liability) of a defendant’s responsibility.
  • Liability apportionment. Once liability is established, this sub-task determines the specific forms of legal responsibility to be imposed, in accordance with civil law. Common forms include monetary compensation, the restitution of damages or cessation of infringement (wan2022civil). Given that monetary compensation is the predominant remedy in tort law (van2001principle), we explicitly add a step of damages calculation where the total loss suffered by the victim is estimated. The total of estimated damages is multiplied by the proportion of liability attributed to each defendant to obtain a precise amount of compensation. Additional remedies, such as restitution or cessation of infringement, are also considered where legally applicable.
@@
Facilitated by the modular and stepwise structure formalized in the LexChain reasoning framework, we introduce a novel task of tort legal reasoning (TLR) to systematically evaluate LLMs’ capacity to perform structured legal reasoning in the context of tort law. We construct an evaluation benchmark, LexChaineval, to assess models’ ability to identify and reason through the critical components involved in tort case analysis. In addition, we propose an LLM-assisted scoring framework to quantify model performance at each reasoning step, enabling interpretable and modular evaluation.
@@
Tort legal reasoning (TLR) requires a model to correctly identify or predict a structured sequence of key legal elements that together form a coherent reasoning chain for tort adjudication. The LexChain framework provides a principled decomposition of the reasoning process into discrete modules, each associated with legally meaningful sub-tasks. We construct the TLR task by closely aligning the task with this formalization, allowing models to be evaluated across the entire reasoning trajectory, from case setup to judgment formulation.
@@
We construct the evaluation dataset LexChaineval based on real-world judicial documents. Authentic judicial decisions offer a realistic, diverse source of tort-related legal scenarios. In particular, cases that proceed to litigation typically involve greater factual and legal complexity, as they often center on issues of conceptual ambiguity or interpretive dispute (schauer2009thinking). These characteristics make litigated cases especially well-suited for modeling and evaluating tort legal reasoning.
We start by sampling 1,000 real-world tort case judgments from cjo2013. To ensure that the evaluation dataset centers on substantive legal reasoning, we exclude cases involving procedural matters, such as voluntary withdrawals, jurisdictional transfers or dismissals on procedural grounds during the sampling process. The resulting dataset comprehensively covers major types of tort disputes, encompassing 45 distinct tort subcategories.
We design an LLM-assisted extraction pipeline to systematically retrieve key legal information from court decision texts. This pipeline extracts essential elements required for tort reasoning, including: plaintiff(s), defendant(s), type of dispute, applicable legal statutes, claims of the plaintiff(s), case facts, total damages, liability determinations and judgment outcomes. For elements commonly expressed in structured formats in court decisions, such as parties, claims, citations of law and judgments, we employ rule-based extraction techniques, including keyword matching and regular expression-based heuristics. For elements requiring holistic understanding and document-wise summarization, such as fact descriptions, liability reasoning and damages calculation, we utilize DeepSeek-V3 (liu2024deepseek_v3) to synthesize relevant information from the court decision texts.
To ensure the quality of data present in LexChaineval, we conduct a rigorous manual validation process on the automatically extracted case information. All extracted data are reviewed and refined by a trained legal annotator, under the supervision of a senior researcher with extensive experience in legal AI and annotation workflows. The extraction and manual validation processes ensure a high-quality benchmark for evaluating tort-related legal reasoning.
@@
To systematically evaluate model outputs on the tort legal reasoning task, we adopt an “LLM-as-a-Judge” evaluation paradigm, widely used in recent literature due to its demonstrated effectiveness and scalability for automated evaluation (zheng2023judging; gu2024survey). We design a structured and automated scoring process using GPT-4o (hurst2024gpt_4o) as the evaluation model. Importantly, instead of relying on subjective, holistic assessments, we decompose the evaluation into a series of discrete, well-defined sub-dimensions and prompt GPT-4o to make rule-based, factual judgments for each sub-dimension. This decomposition mitigates the risk of evaluation bias and enhances reliability of LLM-based evaluation, aligning with existing practices in recent studies (liu2023geval; wang2024dhp; gu2024survey).
The evaluation covers seven distinct scoring dimensions, corresponding directly to the core reasoning elements examined in the LexChaineval benchmark. For each test instance, GPT-4o is instructed to evaluate the model’s response against gold-standard answers, i.e., reference case elements provided in the LexChaineval dataset. The scoring schema for each reasoning element is defined as follows:
@@
    Liability apportionment: Measures the model’s ability to correctly identify the liable parties, the proportion of liability, forms of civil responsibility (e.g., compensation, restitution) and the associated legal justifications.
@@
To assess the robustness and reliability of the LLM-based evaluation, we conduct quality validation where we randomly sample 300 instances from GPT-4o’s ratings and ask two independent annotators to review the ratings. For each of GPT-4o’s ratings, annotators independently assign a binary judgment: a score of 1 if they consider GPT-4o’s evaluation correct, and 0 otherwise. In cases where the two annotators disagree, a senior researcher with expertise in legal AI practices is consulted, and the annotators engage in discussion to reach a consensus. The final agreed-upon labels (0 or 1) are used as the reference for calculating the accuracy of GPT-4o’s ratings. All initial disagreements are also recorded to calculate the inter-annotator agreement score to ensure a fair assessment. Based on this validation criterion, GPT-4o achieves an overall rating accuracy of 0.949 on the sampled set. Inter-annotator agreement, measured using Cohen’s Kappa (cohen1960coefficient), yields a score of 0.619, indicating relatively strong consistency between annotators [^1] According to landis1977measurement, a Cohen’s Kappa score between 0.61 and 0.80 signifies “substantial agreement”..
Constructing LexChain-Style Training Data
@@
To prepare training data enriched with legal reasoning information, we sample approximately 10,600 court decision documents for tort cases. For each case, we first extract key legal elements following the same information extraction practice described in LexChaineval construction. We then leverage DeepSeek-V3 (liu2024deepseek_v3) to assist in the automatic construction of legal reasoning chains, building upon the extracted legal elements. For each step in the reasoning process, we determine the specific types of information required to infer towards the current reasoning objective and construct corresponding reasoning formulas. For example, the identification of litigation parties (i.e., plaintiff(s) and defendant(s)) requires consideration of both the case facts and the claims of the plaintiff(s), while the calculation of compensation details depends on the previously determined liability ratio and the victim’s actual losses. An illustrative example is presented in Figure 2. Each constructed reasoning chain is enclosed in a <<think>> and <</think>> tag pair and prefixed to the final answer consisting of the extracted elements, resulting in the LexChaintrain training data.
@@
We conduct human validation to assess the quality of the reasoning chains automatically generated by DeepSeek-V3. Two annotators independently evaluate 300 samples of LLM-generated reasoning chains based on two assessment criteria: consistency and coherence. Consistency measures the alignment of the generated reasoning text with the original legal elements from which it is constructed, while coherence evaluates whether the output forms a logically sound and well-structured reasoning narrative. The training data constructed by DeepSeek-V3 obtains a consistency rate of 0.973 and a coherence rate of 0.993. The inter-annotator agreement between the two annotators, measured as the Cohen’s Kappa score (cohen1960coefficient), is 0.866, indicating a strong agreement between annotators.
@@
Table 1: Evaluation results on the LexChaineval benchmark. Entries labeled with the model name alone (e.g., GPT-4o) represent the zero-shot inference setting. The variant “w/ PromptLC” denotes models prompted using LexChain-style legal reasoning prompts. Variants “w/ SFTSyll” and “w/ SFTLC” denote models fine-tuned on syllogism-style and LexChain-style SFT data, respectively. The variant “w/ DPOLC” refers to models optimized with LexChain-style DPO data.